OpenStudy (anonymous):

Find equation of the line thats perpendicular to the line 2x +5y+8=0 that passes through the line (-1,-2)

OpenStudy (yttrium):

You know how to find the slope of the given line @jessgrabbinn ???

OpenStudy (anonymous):

y= -2/5x ??

OpenStudy (yttrium):

How'd you get that?

OpenStudy (anonymous):

y= mx+b since the line is 2x+5y+8=0 you would rearrange the formula 5y =-2x -8 y= -2/5x -8/5 am i right?

OpenStudy (yttrium):

Yeah. And perpendicularity means negative reciprocal of the original slope. Then, what is our new slope?

OpenStudy (anonymous):

Hint. Find the slope of the equation 2x +5y+8=0 by transforming it into y=mx + b

OpenStudy (yttrium):

@jessgrabbinn , are you there?

OpenStudy (anonymous):

y= 2/5x + 8/5 ?

OpenStudy (anonymous):

or are you referring to the original line?

OpenStudy (anonymous):

it would be y= -2x/5 - 8/5 so the slope is -2/5 This is the crucial point: Since the equation of the line that you are finding for is perpendicular, the product of their slopes would be equal to -1 or in other words, their slope is negative reciprocal of one another

OpenStudy (yttrium):

Yes. The slope of the perpendicular line is actually the negative reciprocal of the slope of the original line. Get it @jessgrabbinn ??

OpenStudy (anonymous):

Is that my answer? or do I keep going?

OpenStudy (yttrium):

So, the slope of the perpendicular line will be 5/2. And upon using the point slope form. y -y1 = m (x-x1) y +2 = 5/2 (x+1) And there you will find the new line that is perpendicular to the original line. :)

OpenStudy (anonymous):

so the slope of the perpendicular line would be 5/2 then use the point slope form y-y1=m (x-x1) substitute the point (-1,-2) it will become y-(-2) = 5/2 [x-(-1)] simplify: y + 2 = 5/2 (x+1) multiply 2 to both sides 2y + 4 = 5 (x+1) 2y + 4 = 5x + 5 General form of equation is 5x - 2y +1 = 0

OpenStudy (yttrium):

can yuo follow @jessgrabbinn ??

OpenStudy (anonymous):

why did you multiply 2 to both sides?

OpenStudy (anonymous):

@jessgrabbinn that is to remove the denominator in the right side and make the equation simplier

OpenStudy (yttrium):

\[y+2=\frac{ 5 }{ 2 } (x+1)\] to simplify this \[2[y+2=\frac{ 5 }{ 2 } (x+1)]\] So we will have \[2y+4 = 5x + 5\] Hence, the general form will be \[5x-2y+1 = 0\] Are you englihtened now?

OpenStudy (anonymous):

Notice that 5/2 (x+1) became 5 (x+1) because the denominator is cancelled

OpenStudy (anonymous):

OH gotcha, I'm still a little confused on what happened here: " multiply 2 to both sides 2y + 4 = 5 (x+1) 2y + 4 = 5x + 5 General form of equation is 5x - 2y +1 = 0" why is it 5x+5 ?

OpenStudy (anonymous):

how did 5/2x +1 suddenly become 5x+5?

OpenStudy (yttrium):

The 2 is already eliminated in the right side, right? Hence, what is left is 5(x+1) = 5x+5

OpenStudy (anonymous):

where did the 5 come from? after you multiply each side by 2, doesnt it become 2y+4=5x+2 ?

OpenStudy (yttrium):

We actually used distributive property. 5(x+1) = (5x)+(5*1)

OpenStudy (anonymous):

OH

OpenStudy (anonymous):

after you use the distributive prop, shouldnt it become 2y+4=5x+10?

OpenStudy (anonymous):

after you multply by 2

OpenStudy (yttrium):

No. The reason why we multiply it by 2 is that to eliminate the 2 in the denominator. Hence. 5x+5 remains still.